# Quantum Gravity

### v1

@@ 7 Conclusions
We have studied the local limits of nonlocal quantum field theories. Moving along the manifold $\mathcal{M}_{\text{NL}}$ of NL-QFTs by varying a certain parameter $\lambda$, we have inquired whether a target local model is reached when $\lambda$ tends to infinity. When it is so, the nonlocal models can be seen as nonlocal deformations of their target local limits.
The nonlocal deformations are encoded into form factors that multiply the propagators of local theories and remove poles that normally propagate ghosts. We have shown that the form factors inspired by the models mostly studied in the current literature give theories that have well-defined limits only in Euclidean space. Singular behaviors appear in the Minkowskian correlation functions when $\lambda\rightarrow\infty$.
To overcome this difficulty, we have relaxed certain requirements and defined a new class $\mathcal{M}_{\text{AL}}$ of unitary, asymptotically local theories, so-called because they have well-defined local limits in Minkowski spacetime. Unitarity forbids target models with ghosts and privileges models with purely virtual particles.
Inside the bubble diagram, the nonlocal deformation generates PVPs directly, in the local limit. In the triangle diagram, it does so possibly up to multi-threshold corrections, which can be adjusted by tuning the deformation itself.
The asymptotically local deformation of a local theory is not unique, and not renormalizable. In our approach, this is not a liability, because we are not proposing AL-QFTs as candidate fundamental theories of nature, but merely as tools to provide alternative formulations of theories with PVPs, approximations to study the violations of microcausality and the peak uncertainties, or alternative approaches to off-shell physics, more commonly treated by restricting QFT to a finite interval of time and a compact space manifold. In that case, the non uniqueness is associated with the possibility of introducing parameters that describe the apparatus, the classical environment surrounding the experiment, and the observer itself.
When this type of arbitrariness is present, the first thing to do is identify the parameters that are relevant to the phenomenon under observation. If they are finitely many, as is reasonable to expect in normal circumstances (or cleverly arranged setups), one sacrifices an equal number of initial measurements to fix them. At that point, one can verify whether the subsequent measurements are correctly predicted or not.
@@ Acknowledgments
A Bubble diagram with PVPs
We review the calculation of the bubble diagram with one or two circulating PVPs, in the formulation of ref. [12]. We work in Minkowski spacetime, so here $p$ and $k$ denote Minkowskian momenta.
@@ A Bubble diagram with PVPs
for each term inside the parentheses, where $\mathcal{P}$ is the Cauchy principal value. We obtain
$\mathcal{B}_{\text{ph}}$ $=$ $\mathcal{P}\int\frac{\mathrm{d}^{D-1}\mathbf{p}}{(2\pi)^{D-1}}\frac{i}{4\omega_{\mathbf{p}}\omega_{\mathbf{p}-\mathbf{k}}}\left(\frac{1}{k^{0}+\omega_{\mathbf{p}-\mathbf{k}}+\omega_{\mathbf{p}}}-\frac{1}{k^{0}-\omega_{\mathbf{p}-\mathbf{k}}-\omega_{\mathbf{p}}}\right)$ (A.1)
@@
$\mathcal{B}_{\text{PVP}}=\mathcal{P}\int\frac{\mathrm{d}^{D-1}\mathbf{p}}{(2\pi)^{D-1}}\frac{i}{4\omega_{\mathbf{p}}\omega_{\mathbf{p}-\mathbf{k}}}\left(\frac{1}{k^{0}+\omega_{\mathbf{p}-\mathbf{k}}+\omega_{\mathbf{p}}}-\frac{1}{k^{0}-\omega_{\mathbf{p}-\mathbf{k}}-\omega_{\mathbf{p}}}\right).$ (A.2)
For the generalization to the other diagrams, see [12].

### v3

@@ 6 Conclusions
We have investigated the relationship between nonlocal and local quantum field theories, and searched for a practical definition of local limit. On general grounds, it is not obvious that a nonlocal model admits a local limit of some sort.
The nonlocal deformations are encoded into form factors that multiply the propagators and remove poles that otherwise would propagate ghosts. We have shown that the form factors inspired by the models studied in the current literature give theories that have well-defined limits only in Euclidean space. Singular behaviors appear in the Minkowskian correlation functions.
To overcome this difficulty, we have relaxed certain requirements and defined a new class of unitary, asymptotically local theories, which have well-defined local limits in Minkowski spacetime. Unitarity forbids target models with ghosts and privileges models with purely virtual particles.
Inside the bubble diagram, the nonlocal deformation generates PVPs directly in the limit. In the triangle diagram, it does so possibly up to multi-threshold corrections, which can be adjusted by fine tuning the deformation itself.
The asymptotically local deformation of a local theory is not unique, and not renormalizable. This is not a liability, in our approach, because we are not proposing AL-QFTs as candidate fundamental theories of nature, but merely as tools to provide alternative formulations of theories with PVPs, approximations to study the violation of microcausality and the peak uncertainty, or alternative approaches to off-shell physics, more commonly treated by restricting QFT to a finite interval of time and a compact space manifold.
In off-shell physics, the non uniqueness is associated with the possibility of introducing parameters that describe the apparatus, the classical environment surrounding the experiment and the observer itself. At the practical level, the theory is predictive in the following sense. One first identifies the parameters that matter to the phenomenon under consideration. Assuming that they are finitely many– as is reasonable to expect under normal circumstances, or in cleverly arranged setups –, one sacrifices an equal number of initial measurements to fix them. After that, one can verify whether the subsequent measurements are correctly predicted or not.
We believe that the nonlocality-locality relation worked out here can stimulate further investigations and some out-of-the-box thinking.
@@ Acknowledgments
A Key functions
In this appendix we give entire functions that tend to the absolute value, the sign function and the Cauchy principal value in the local limit $\lambda\rightarrow\infty$. They are used in section 2 to build the models studied in the paper.
A.1 Approximating the absolute value
$h(z)\equiv\exp\left(\frac{1}{2}\int_{0}^{z^{2}}\frac{1-\mathrm{e}^{-w}}{w}\mathrm{d}w-\frac{\gamma_{\text{E}}}{2}\right)=\exp\left(\frac{1}{2}\ln z^{2}+\frac{1}{2}\Gamma(0,z^{2})\right).$ (A.1)
Precisely, in the double cone $\mathcal{C}=\{z:-\pi/4<$ arg[$z$] $\pi/4$ or $3\pi/4<$ arg[$z$] $5\pi/4\}$, we have [26]
$h(z)=\sqrt{z^{2}}\left[1+\frac{\mathrm{e}^{-z^{2}}}{2z^{2}}\left(1+\mathcal{O}\left(\frac{\mathrm{e}^{-z^{2}}}{z^{2}}\right)\right)\left(1+\mathcal{O}\left(\frac{1}{z^{2}}\right)\right)\right],$ (A.2)
so
$\lim_{\lambda\rightarrow+\infty}\frac{h(\lambda z)}{\lambda}=\sqrt{z^{2}}\text{\quad for }z\in\mathcal{C},\qquad\lim_{\lambda\rightarrow+\infty}\frac{h(\lambda x)}{\lambda}=|x|\quad\text{\ for }x\in\mathbb{R}.$ (A.3)
It is important to stress that $h(z)$ has no zeros. Its reciprocal $1/h(z)$ is useful to build the “propagators” that tend to purely virtual particles in the local limit.
The function $h(z)$ is even,
h(z)=h(z),h(z)=h(-z), (A.4)
and satisfies
$h^{\ast}(z)=h(z^{\ast}).$ (A.5)
Moreover, on the real axis it is positive and bounded from below by the absolute value, as illustrated in fig. 2:
$h(x)\geqslant|x|,\qquad x\in\mathbb{R}\text{.}$ (A.6)
Finally, (A.3) gives
$\lim_{\lambda\rightarrow+\infty}\frac{h^{2}(\lambda z)}{\lambda^{2}z}=z\quad\text{for }z\in\mathcal{C}.$ (A.7)
Figure 2: Plot of $h(x)$ for real $x$
A.2 Approximating the sign function
The sign function sgn(x)(x) can be approximated on the complex plane by the entire function
$\sigma(z)=\frac{z}{h(z)},$ (A.8)
$\lim_{\lambda\rightarrow+\infty}\sigma(\lambda z)=\text{sgn}(\text{Re}[z])\quad\text{for }z\in\mathcal{C},\qquad\lim_{\lambda\rightarrow+\infty}\sigma(\lambda x)=\text{sgn}(x)\quad\text{for }x\in\mathbb{R}.$ (A.9)
Note that $\sigma(z)$ is odd in the complex plane:
$\sigma(z)=-\sigma(-z).$ (A.10)
Moreover, (A.6) gives
$|\sigma(x)|\leqslant 1,\qquad x\in\mathbb{R}\text{.}$ (A.11)
A.3 Approximating the principal value
The combination $\sigma(x)/h(x)=x/h^{2}(x)$ approximates the Cauchy principal value $\mathcal{P}$ on the real axis. Specifically, we prove that
$\lim_{\lambda\rightarrow+\infty}\frac{\lambda\sigma(\lambda x)}{h(\lambda x)}=\mathcal{P}\frac{1}{x},\qquad x\in\mathbb{R},$ (A.12)
in the sense of distributions.
First observe that
$\lim_{\lambda\rightarrow+\infty}\frac{\lambda\sigma(\lambda x)}{h(\lambda x)}=\frac{1}{x}$ (A.13)
for every real $x\neq 0$. This result follows from (A.3) and (A.9).
If $\varphi(x)$ denotes a real test function, consider the integral
$\mathcal{I}\equiv\lim_{\lambda\rightarrow+\infty}\int_{-\infty}^{+\infty}\mathrm{d}x\frac{\lambda\sigma(\lambda x)\varphi(x)}{h(\lambda x)},$ (A.14)
which can also be written as
thanks to (A.4) and (A.10). The modulus of the $\mathcal{I}$ integrand is bounded above by a $\lambda$-independent function that is integrable on $\mathbb{R}$. Indeed, the inequalities (A.6) and (A.11) give
$\left|\frac{\lambda\sigma(\lambda x)(\varphi(x)-\varphi(-x))}{2h(\lambda x)}\right|\leqslant\frac{1}{2}\left|\frac{\varphi(x)-\varphi(-x)}{x}\right|,$ (A.15)
for real $x$. Then the dominated convergence theorem allows us to exchange the limit with the integral. Given that (A.13) holds almost everywhere, we obtain
$\mathcal{I}=\int_{-\infty}^{+\infty}\mathrm{d}x\frac{\varphi(x)-\varphi(-x)}{2x}=\mathcal{P}\int_{-\infty}^{+\infty}\frac{\mathrm{d}x}{x}\varphi(x),$ (A.16)
as we wished to show.
B Bubble diagram with PVPs
In this appendix we review the calculation of the bubble diagram with one or two circulating PVPs, in the formulation of ref. [11]. We work in Minkowski spacetime, so here $p$ and $k$ denote Minkowskian momenta.
@@ B Bubble diagram with PVPs
for each term inside the parentheses. We obtain
$\mathcal{B}_{\text{ph}}$ $=$ $\mathcal{P}\int\frac{\mathrm{d}^{D-1}\mathbf{p}}{(2\pi)^{D-1}}\frac{i}{4\omega_{\mathbf{p}}\omega_{\mathbf{p}-\mathbf{k}}}\left(\frac{1}{k^{0}+\omega_{\mathbf{p}-\mathbf{k}}+\omega_{\mathbf{p}}}-\frac{1}{k^{0}-\omega_{\mathbf{p}-\mathbf{k}}-\omega_{\mathbf{p}}}\right)$ (B.1)
@@
$\mathcal{B}_{\text{PVP}}=\mathcal{P}\int\frac{\mathrm{d}^{D-1}\mathbf{p}}{(2\pi)^{D-1}}\frac{i}{4\omega_{\mathbf{p}}\omega_{\mathbf{p}-\mathbf{k}}}\left(\frac{1}{k^{0}+\omega_{\mathbf{p}-\mathbf{k}}+\omega_{\mathbf{p}}}-\frac{1}{k^{0}-\omega_{\mathbf{p}-\mathbf{k}}-\omega_{\mathbf{p}}}\right).$ (B.2)
For calculations of more complicated diagrams and the general proof of unitarity, see [11].
It is important to stress that if we adopt the principal value we see in (4.7) as an alternative propagator in Feynman diagrams, we obtain a theory of “Wheelerons” [27], which propagates ghosts. For example, the bubble diagram gives
$\mathcal{P}\int\frac{\mathrm{d}^{D}p}{(2\pi)^{D}}\frac{1}{p^{2}-m^{2}}\frac{1}{(p-k)^{2}-m^{2}},$ (B.3)
which has a nonvanishing, unphysical absorptive part [28], absent in (B.2). The moral of the story is that in a theory of PVPs the radiative corrections are not given by the usual diagrams with a different propagator, but by new combinations of diagrams [12, 11].